For the reaction PCl3(g)+Cl2(g)=PCl5(g),the value of Kc at 523K is 26.Then value of Kp at this temperature is(a) 0.61 (b) 0.57(c) 0.83 (d) 0.46

Grade:12th pass

1 Answers

sneha
6 years ago
kp/kc = (RT)^∂n
kp= kc*(RT)^∂n
kc= 26
R=0.082
kp= 26*(0.082*523)^(-1)
    =26*0.023
    =0.606 =0.61
 
Ans :- a
i think this is the answer. if it is wrong plz inform me and give the right answer... thanks...i dont know is it correct..iam preparing for neet for next year. okay have a nice day.